Complete the lab 5SERIES CIRCUITS VW REFERENCE READING Principles of Electric Circuits: Sectang 81 through 6-4. RELATED PROBLEMS FROM PRINCIPLEPROCEDURE 1. Calculate the upper and lower limits of the resistance for each resistor and re- cord these in Table 8-1. 2. UseDATA FOR EXPERIMENT 8 TABLE 8-1 1k02 2k0 3.9 k125.1 km Rx = 15 kn Rule Rau Rr TABLE 8-2 Inom (mA) (mA) ܩܢܕܐ BLE V, - 10 V 7 IName Date QUESTIONS FOR EXPERIMENT 8 1. If the moagured value of Ry did not equal the sum of the indi- vidually measured resi
Answer #1

Currently iam in home . I don't have any equipment like DMM. So, I am calculating manually.

=> Given circuit is, I aimat {R3 Ry Table 8-1:- 1) As per question, they give (+5%) tolerance 282 јко 13k2 3.9612 . 5.1k2 | R2) 73) FIESTAS I dont have any instrument to measure resistance. from Table; a Rt (min) - 14.25 km2 = 095+19K 72-85k+370.5k7 Table 8-3: - | Unom (v) | 7.5* | I=0.5mA | Vmen cu) | 7.125 | Vmax(u) 7.875 CY . 7 I mean can be measured using Dmm. we meaIQ YR 3) >> To calculate Imrin we need to 28 maximum resistance So, given tolerance +50% >> Because For Iks, akr and 3k2 JI3) The cussent through each asistov in a sexies-connected (percuit such as that in figure 8-1 is the same I Sol True In a ser5). If any resistance value reduced in series circuit, the total resistance will also reduced, then automatically current will increase.

Because , current and resistance are inversely proportional. So, minimum resistance will lead to maximum current.